# Unfolding the Literature: A Review of Robotic Cloth Manipulation

### v1

@@ 7.2 Household chores
7.3 Fashion Industry
The fashion industry presents numerous opportunities for autonomous robots with adaptive capabilities. This subsection focuses on specific applications such as manufacturing, dyeing, and recycling, directly linked to adaptability in handling, perceiving, and quality management of textiles. Given the extensive previous discussion on sorting, smoothing, ironing, and folding, this section will concentrate on these industry-specific applications.
In cloth manifacturing, robots are increasingly integrated into the cutting and sewing stages of garment production [141], with potential advancements enabling them to detect variations in fabric properties like thickness and stretchability. This ability could allow precisely adjust techniques for each material, significantly reducing waste and enhancing resource efficiency, thus improving sustainability and garment quality.
@@ 7.3 Fashion Industry
Finally, recycling is crucial for sustainability in the fashion industry [139]. Similar to sorting, enhanced interactive perception of fabric types, construction methods, and material conditions can improve this process by accurately identifying textiles suitable for recycling [70].
@@ 8 DISCUSSION AND FUTURE PERSPECTIVES
This section summarizes the main concepts discussed in this review, further identifying open problems and detailing future research directions and grand challenges to foster the development of perception and manipulation skills that generalize to variations of textile properties and manipulation tasks.
8.1 Synergies
In previous sections, we examined modeling, perception and manipulation separately. However, there is significant interplay among these domains, impacting generalization and adaptation capabilities. Here, we review some challenges faced by these components individually and how their integration can be beneficial for generalization.
Offloading most computational effort to modeling, such as using analytical models, allows for offline optimization of manipulation trajectories. This reduces the burden on control strategies and perception, as optimal strategies can faithfully execute the task in an open loop [21]. While cloth models facilitate transferring to novel tasks [84], real-world applications require perception techniques to adjust simulator parameters for each real-world object as in Section 3.3.2. This involves estimating object properties from sensors or feedback and replanning to address inaccuracies, making perception crucial for generalization [82, 48].
Model-free learning techniques such as RL and IL shift the computational complexity to manipulation by learning end-to-end from raw data [88]. While this allows the reduction of the requirements from the perception module, it has to trade sample efficiency and generalizability, which often results in training a single policy for each task, limiting flexibility in varying environments. However, not all the components of a sequential manipulation task need to be learned end-to-end and offloaded to the policy. Sub-tasks within cloth manipulation can be simplified with perception modules. Examples include flattening based on wrinkle [27] or grasp point detection [64], and folding plans based on corner alignment, which delegate specific tasks to perception, bypassing the need for extensive learning.
Conversely, relying solely on perception and heuristic-based approaches hinders generalizability, as it is challenging to design heuristics for all real-world scenarios. Learning policies can address tasks where heuristic-based algorithms fall short. For example, which point to grasp could be addressed by the perception algorithm, how to reach the grasping point can be resolved with a standard planning algorithm, and how to optimize the manipulation once the cloth is grasped can be learned and performed by an end-to-end policy. The underexplored area of switching between learning and heuristic approaches and determining when to use each presents a promising research direction. This is particularly relevant with the advent of foundation models capable of reasoning about semantics and tasks and adapting to rules and human preferences.
@@ 8.2 Open Problems
To discuss the generalization and adaptation of current methods for cloth manipulation, we provided an overview of fundamental textile properties that contribute to the variations of real-world textiles, highlighting the complexity of understanding their role in manipulation due to the intertwined role of each of them. However, the influence of each textile property in robotic manipulation is still underexplored. Moreover, the extent to which explicit identification of each property is required remains unclear. An open avenue for research is identifying a subset of pertinent features for different manipulation task, which can be used to determine the extent to which the textile properties need to be identified.
@@
Adding to the complexity of identifying and defining the textile properties is the dynamic nature of these properties under varying conditions. Aging of the cloth, as well as wet, dirty, or dry conditions, can alter a textile’s behavior, emphasizing the need for continuous sensing and adaptive perception systems in autonomous agents. Moreover, physical attributes like thickness, softness, and durability [72] are typically evaluated alongside semantic descriptors, like smooth, absorbent, hairy, and slippery that humans use to characterize textiles [81]. Particularly, with the advances in Large Language Models (LLMs) there is an emerging potential to bridge these descriptive terms with a physical understanding of textile properties [145].
@@
One key aspect for manipulating a variety of textiles and addressing the variation of properties is perception. While interactive perception with different types of sensors and exploratory actions can enhance adaptability and reduce uncertainty in presence of partial observability, it remains underexplored. Multimodal sensing emerges as a pivotal strategy in this context, integrating various sensory inputs like tactile, visual, and auditory data to provide a more holistic understanding of textile properties. This approach holds significant potential, as it allows robots to leverage multiple temporal information sources, compensating for the limitations of individual sensors.
@@
A notable finding from our review of manipulation techniques is the limited amount of works demonstrating effective generalization to a wide range of deformable objects, with only a few exploring adaptation techniques. Most current efforts utilize domain randomization in simulations to enhance sim2real transfer. While promising, these methods often overlook the potential of various domain adaptation techniques that could allow robots to dynamically adjust their manipulation strategies based on real-time environmental feedback and changes in textile properties. Similarly, the use of state-of-the-art models such as diffusion models, in combination with imitation learning techniques, has predominantly focused on rigid objects due to the complexity of state estimation and high-dimensional state spaces of textiles. This leaves open questions about their applicability to deformable objects and the challenges of deploying these methods without assuming rigidity.
@@
Building on the necessity for improved generalization in manipulation techniques, benchmarking, standardization and datasets emerge as essential yet challenging areas. The diversity of robotic embodiments, sensor configurations, and test sets, complicates the comprehensive comparison of approaches and algorithms. Specifically, the development of universally applicable testing datasets is hindered by issues like stock availability and the difficulties associated with measuring properties accurately. Although there has been some progress towards creating standardized test sets and benchmarks, these continue to be significant open problems within the field. Additionally, a major challenges in dataset availability and scalability is accurately labeling the deformable object dataset, which limits their widespread distribution and application.
@@ 8.3 Grand Challenges
To push the boundaries of robotic manipulation of textiles, we identify the following critical challenges that the research community should address:
Perception of properties of novel objects: Develop advanced methods for accurately estimating key physical and semantic properties of new textiles. This challenge requires integrating sophisticated sensing technologies and exploratory procedures to identify these properties effectively. Achieving this would enhance robotic manipulation, enabling robots to adaptively handle a diverse range of textiles.
Multi-modal, Multi-task, Multi-embodiment Generalist Agent: Develop an agent capable of addressing multiple tasks in unknown environments, especially those involving the manipulation of deformable objects and assistive tasks with humans in the loop. This requires the agent to autonomously adapt to complex situations and acquire skills from minimal human demonstrations, leveraging multi-modal sensory inputs and versatile embodiment strategies. Achieving this would significantly advance robotic flexibility and effectiveness in diverse real-world applications.
Datasets and Benchmarking: Address the challenges in creating comprehensive datasets for high-dimensional state spaces and degrees of freedom in robotic manipulation. Develop a well-annotated dataset that captures 2D or 3D states across diverse, dynamic interactions labeled with relevant features and properties. Additionally, design standardized experimental benchmarking procedures for textile objects, considering stock availability and object variations, to facilitate consistent progress across research institutions. Achieving this will significantly enhance the community’s ability to develop and compare advanced robotic systems.

### v2

@@ 7.2 Household chores
7.3 Textile Industry
The textile industry, encompassing sectors such as fashion, automotive, and construction, presents numerous opportunities for autonomous robots with adaptive capabilities. This subsection focuses on specific applications such as manufacturing, dyeing, and recycling, directly linked to adaptability in handling, perceiving, and quality management of textiles. Given the extensive previous discussion on sorting, smoothing, ironing, and folding, this section will concentrate on these industry-specific applications.
In cloth manufacturing, robots are increasingly integrated into the cutting and sewing stages of garment production [145], with potential advancements enabling them to detect variations in fabric properties like thickness and stretchability. This ability could allow precisely adjusting techniques for each material, significantly reducing waste and enhancing resource efficiency, thus improving sustainability and garment quality.
@@ 7.3 Textile Industry
Finally, recycling is crucial for sustainability [143]. Similar to sorting, enhanced interactive perception of fabric types, construction methods, and material conditions can improve this process by accurately identifying textiles suitable for recycling [71].
@@ 8 DISCUSSION AND FUTURE PERSPECTIVES
In previous sections, we examined modeling, perception and manipulation separately. However, there is significant interplay among these domains, impacting generalization and adaptation capabilities. Offloading computational effort to modeling allows for model-based optimization of manipulation trajectories, reducing the burden on control and perception [21]. However, real-world applications need perception techniques to align the parameters of the model to the real-world object, making perception crucial for generalization [48, 83]. While model-free learning techniques like RL and IL learn end-to-end from raw data [89], reducing perception needs, they sacrifice sample efficiency and generalizability to novel variations of the environment and tasks. Nonetheless, not all components of manipulation tasks need end-to-end learning; perception modules can simplify sub-tasks like flattening wrinkles [27], grasp point detection [65], and folding plans, bypassing the need for extensive learning. As an example, perception can determine grasp points, how to reach the grasping point can be resolved with a standard planning algorithm, and how to optimize the manipulation once the cloth is grasped can be learned and performed by an end-to-end policy. Exploring when to switch between learning and heuristics is a promising research direction. This is particularly relevant with the advent of foundation models capable of reasoning about semantics, sequential tasks, and adapting to rules and human preferences.
In the remainder of the section, we further identify open problems and detail future research directions and grand challenges to foster the development of perception and manipulation skills that generalize to variations of textile properties and manipulation tasks.
@@ 8.1 Open Problems
To discuss generalization and adaptation in cloth manipulation, we reviewed fundamental textile properties and their complex, intertwined roles. However, the influence of each property in robotic manipulation is still underexplored, and the necessity of explicitly identifying each property remains unclear. An open avenue for research is identifying a subset of pertinent features for different manipulation tasks to determine the extent to which textile properties need to be identified.
@@
Adding to the complexity of defining textile properties is their dynamic nature under varying conditions. Aging, wetness, dirt, and dryness can alter a textile’s behavior, necessitating continuous sensing and adaptive perception in autonomous agents. Physical attributes like thickness, softness, and durability [73] are often evaluated alongside semantic descriptors like smooth, absorbent, hairy, and slippery [82]. Advances in Large Language Models (LLMs) present potential to bridge these descriptors with a physical understanding of textile properties [149].
@@
One key aspect for manipulating a variety of textiles and addressing the variation of properties is perception. While interactive perception with different types of sensors and exploratory actions can enhance adaptability and reduce uncertainty, it remains underexplored. Multimodal sensing emerges as a pivotal strategy in this context, integrating various sensory inputs like tactile, visual, and auditory data to provide a more holistic understanding of textile properties. This approach holds significant potential, as it allows robots to leverage multiple temporal information sources, compensating for the limitations of individual sensors.
@@
A significant observation of our review of manipulation techniques is the limited amount of work demonstrating effective generalization to a wide range of deformable objects, with few exploring adaptation methods. Most efforts use domain randomization in simulations for sim2real transfer, often overlooking domain adaptation techniques that allow dynamic adjustment based on real-time feedback due to perceptual challenges. Similarly, state-of-the-art models, such as diffusion models, in combination with imitation learning techniques, remain underexplored. This leaves open questions about their applicability to deformable objects and the challenges that might arise due to the complexity of state estimation and high-dimensional state spaces of textiles.
@@
Building on the necessity for improved generalization in manipulation techniques, benchmarking, standardization, and datasets emerge as essential yet challenging areas. The diversity in robotic embodiments, sensor configurations, and test sets complicates comprehensive comparisons. Developing universally applicable testing datasets faces challenges like stock availability and accurate property measurement. Although progress has been made, standardizing test sets and benchmarks remains a significant problem. Additionally, a major challenge in dataset availability and scalability is accurately labeling deformable object datasets, limiting their widespread distribution and application.
@@ 8.2 Grand Challenges
To push the boundaries of robotic manipulation of textiles, we identify the following critical challenges that may have major breakthroughs in the coming years:
(i) Perception of properties of novel objects for accurately estimating key physical and semantic properties of new textiles, enabling robots to reduce the uncertainties about the environment and adaptively handle a diverse range of textiles.
(ii) Adaptive multi-task and multi-modal agent capable of autonomously adapting to complex situations, performing long-horizon tasks, and using multi-modal sensory inputs to navigate uncertainties, enabling seamless integration into homes, industries, and healthcare facilities.
(iii) Novel datasets and benchmarks capturing real-world variations in object properties, physical interactions, and manipulation tasks to facilitate standardized benchmarking and enable consistent comparisons of robotic systems across research institutions.
The field of robotic manipulation of textiles is extensive and includes numerous important technological areas not mentioned here. Thus, the list provided is not exhaustive. The three grand challenges identified — perception of properties of novel objects, adaptive multi-task and multi-modal agents, and novel datasets and benchmarks — represent critical areas that have the potential to drive major advancements. These challenges encompass core perceptual technologies, general and adaptive capabilities, and standardized evaluation methods, aiming to enhance the flexibility and effectiveness of robotic systems in diverse real-world applications.
AUTHORS CONTRIBUTIONS
A.L. conceived the idea and the structure of the review, authored sections 1,2,4,8, coauthored sections 3,5, revised the full manuscript, created Figures 1 and 3, contributed to the creation of Tables 1 and 2, and coordinated the overall preparation and revision of the manuscript. Y.W. authored section 5, revised the full manuscript, and contributed to the creation of Table 1. I.G.-C. authored section 6, coauthored section 2, revised the full manuscript, and created Figure 2. D.B.-M. coauthored section 5, revised the full manuscript, and contributed to the creation of Table 1. M.M. authored section 7, revised the full manuscript, and contributed to the creation of Table 2. M.W. assisted with the delineation of the structure of the review, contributed to section 3, and revised the full manuscript. H.Y. coauthored section 3 and revised the full manuscript. Z.E. coauthored section 7 and revised the full manuscript. D.H. coauthored section 5 and revised the full manuscript. J.B. coauthored section 6 and revised the full manuscript. G.A. and D.K. revised the entire manuscript.
ACKNOWLEDGMENT
This work has been supported by the European Research Council (ERC-BIRD); the European Union’s Horizon Europe Programme through projects SoftEnable (HORIZON-CL4-2021-DIGITAL-EMERGING-01-101070600) and IRE (HORIZON-CL4-2023-DIGITAL-EMERGING-01-101135082); the Swedish Research Council, the Knut and Alice Wallenberg Foundation, and the National Science Foundation under NSF CAREER grant number IIS-2046491; the project ROB-IN PLEC2021-007859 funded by MCIN/ AEI /10.13039/501100011033 and by the European Union NextGenerationEU/PRTR.